Troll Island is a small island nation that recently experienced an autonomous change in aggregate expenditures (AE). AE increased by 3 billion, and the marginal propensity to consume on Troll Island is equal to 0.65 . What is the change in Troll Island's real GDP after the increase in AE? Enter your answer in billions of dollars, rounded to one place after the decimal. For example, an answer of $2,500,000 should be entered as 2.5.
Answer #1

Answer

Multiplier =1/(1-MPC)

MPC=marginal propensity to consume=0.65

Multiplier =1/(1-0.65)

=2.85714286

change in GDP = change in AE *Multiplier

=3*2.85714286

=8.57142858

=8.6 billion

the change in Troll Island's real GDP after the increase in AE is $8.6 billion